40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1420     Entry time: Tue Mar 24 09:04:02 2009
Author: steve 
Type: Update 
Category: SUS 
Subject: 4.8 mag earthquake 

SRM, ITMX, ETMX, ITMY and ETMY lost damping at 4:55am this morning from 4.8 magnitude earthquake.

Their damping were restored.

C1:SUS-ITMX_URSEN_OUTPUT swich was found in off position. It was turned on.

MZehnder  and MC were locked.

The WFS qpd spot needs recentering

ELOG V3.1.3-